OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (8, 12) and has a slope of −2? y = x +

OpenStudy (alexandervonhumboldt2):

well first is y=-2x+b to find b substitute the coordinates into the equation.

OpenStudy (alexandervonhumboldt2):

12=-2*8+b b=28 y=-2x+28

OpenStudy (alexandervonhumboldt2):

slope is the cooeficent with the x

OpenStudy (anonymous):

thxs ca u help me with some more please?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

also welcome to openstudy!!!

OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (3, 1) and (2, 9)? y = x +

OpenStudy (anonymous):

oh thxs:)

OpenStudy (alexandervonhumboldt2):

slope formula is (y1-y2)/(x1-x2)=(9-1)/(2-3)=-8 slope is -8. y=-8x+b 1=-8*3+b b=25---> y=-8x+25

OpenStudy (anonymous):

so this the y=-8x+25 answer:)?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):